# Neuronal correlations shape the scaling behavior of memory capacity and nonlinear computational capability of recurrent neural networks

### v2

@@ VII Discussion
In the present study, we have investigated the relationship between the number of readout neurons and the computational capacity for large random RNN reservoirs. Through analytical and numerical approaches, we have demonstrated that the memory capacity grows sublinearly with $L$, and that nonlinear computational capabilities emerge incrementally with $L$, enabling the reservoir to handle increasingly complex tasks. The specific memory and nonlinearity requirements of a given task can be assessed using the framework proposed by [23]. In this context, our results provide a practical principle for designing cost-effective RC systems: while increasing $L$ raises computational and structural costs and yields diminishing returns for memory, it becomes essential when tasks demand higher-order nonlinear computations.
@@
Our findings have implications that extend beyond the framework of RC. First, in neuroscience, the impacts of neuronal correlations on function have been extensively investigated from the perspective of neural manifold [24, 25, 26]. A neuronal network with a high level of correlations occupies a restricted region of $N$-dimensional phase space, whose linear dimension is quantified by effective dimension [27]. In the context of population coding, a well-established trade-off exists between representation richness and robustness: high-dimensional (low-correlation) activity enables complex and diverse information encoding, while low-dimensional (high-correlation) activity supports redundant encoding and noise robustness [25, 26]. However, the mammalian cortex in the awake state is known to operate in an asynchronous regime characterized by near-zero neuronal correlations [28, 29, 30], akin to those observed in our large random RNNs. In addition, these prior studies have only focused on the implications of neuronal correlations for static information processing, i.e., mapping static inputs to outputs. Our findings suggest a novel role for weak correlations in dynamic information processing, offering a new hypothesis about their functional significance in the brain.
@@ Appendix A THEORETICAL ANALYSIS ON STATISTICS OF NEURONAL ACTIVITY
In this section, we derive statistics of neuronal activity based on dynamical cavity approach [20, 36]. Several results and methodologies introduced here are frequently utilized in deriving memory capacity in Appendix.B.
In deriving single neuronal statistics such as the variance of $x_{i}(t)$ (see Appendix.A.1) in the limit of $N\to\infty$, the same result can be obtained via dynamical mean-field theory [12, 19] . However, since the dynamical mean-field theory transforms the $N$-body system into the effective single-body system, it is inherently unable to evaluate many-body neuronal statistics, such as neuronal correlations (see Appendix.A.2) and memory capacity(see Appendix.B.2). In contrast, the dynamical cavity method is, in principle, applicable to $n$-body neuronal statistics for any $n$. Furthermore, while the aforementioned limitation of dynamical mean-field theory has recently been addressed through subleading corrections to the saddle-point solution [37, 38], the dynamical cavity method remains advantageous for its conceptual clarity.
@@
Before delving into the details, we first outline the necessity of employing the dynamical cavity method. Consider calculating the quenched variance of neuronal activity, $[x_{i,t}^{2}]$. For simplicity, we ignore input and noise signals. According to time evolution equation of neurons (Eq.(1)), $x_{i,t}$ is the sum of a large number ($N\gg 1$) of mutually independent quantities, $J_{ij}\phi(x_{j,t-1})$. This implies that $x_{i,t}$ follows a Gaussian distribution due to the central limit theorem, and thus, it is sufficient to determine the complete statistics of $x_{i,t}$ by calculating its mean and variance. Given the symmetry of the model, we have $[x_{i,t}^{2}]=0$. On the other hand, we can intuitively calculate $[x_{i,t}^{2}]$ as:
$[x_{i,t}^{2}]=\sum_{j,k}^{N}\left[J_{ij}J_{ik}\phi_{j,t-1}\phi_{k,t-1}\right]\overset{?}{=}\sum_{j,k}^{N}\left[J_{ij}J_{ik}\right]\left[\phi_{j,t-1}\phi_{k,t-1}\right]=g^{2}\left[\phi_{j,t-1}^{2}\right]=g^{2}\langle\phi(x)\rangle_{x\sim\mathcal{N}(0,[x_{i,t-1}^{2}])},$ (13)
yielding time series of $[x_{i,t}^{2}]$. Here we assumed in the second equation that the recurrent weights, $J_{ij}$ and $J_{ik}$, are independent of the neuronal activities, $\phi_{j}$ and $\phi_{k}$. Fortunately, this assumption is validated in the limit of $N\to\infty$, as shown with the dynamical mean-field theory. However, this assumption is likely to break down when finite size effects are taken into account. To circumvent this challenge, we need to employ the dynamical cavity method.
@@ A.1 Single neuronal statistics
We aim to derive the quenched average of the time-variance of neuronal activity, $[\langle x_{i}^{2}\rangle]$. Note that due to the self-averaging, it is equivalent to the population average of time-variance to leading order, i.e. $[\langle x_{i}^{2}\rangle]=\frac{1}{N}\sum_{i}^{N}\langle x_{i}^{2}\rangle$. The symmetry of the model setting ensures that the quenched average of the time-mean, $[\langle x_{i}\rangle]$, vanishes.
The overview of dynamical cavity method is described in Fig.6. One auxiliary neuron, indexed by 00, is added to the original reservoir RNN whose neurons are indexed by $i=1,\cdots,N$. The original neuron is perturbated by the addition of the auxiliary neuron. By denoting the perturbation applied to the $i$th neuron at time $t$ by $\delta_{i,t}$, we describe the dynamics of the auxiliary neuron as
@@
We define $\chi_{ij,ts}\equiv\delta x_{i,t}/\delta I_{j,s}$ as the linear susceptibility of the $i$th neuron at time t, induced by an infinitesimal external inputs applied to the $j$th neuron at time $s\leq t$, $I_{j,s}$. Then, as the perturbation to $i$th neuron at time $t$, $\delta_{i,t}$, is the sum of all direct perturbations to $j\in\{1,\cdots,N\}$ at time $s<t$ weighted by the corresponding susceptibility $\chi_{ij,ts}$, we obtain
$\delta_{i,t}=\sum_{s<t}\sum_{j=1}^{N}\chi_{ij,ts}J_{j0}\phi_{0,s}.$ (15)
@@
$x_{0,t+1}$ $=\underbrace{\sum_{i=1}^{N}J_{0i}\phi_{i,t}}_{\equiv\eta_{t}}+\sum_{s<t}\underbrace{\sum_{i,j}^{N}\chi_{ij,ts}J_{0i}J_{j0}\phi^{\prime}_{i,t}}_{\equiv\kappa_{s}}\phi_{0,s}+J_{00}\phi_{0,t}+u_{0}s_{t+1}+\xi_{0,t+1}$
$\equiv\eta_{t}+\sum_{s<t}\kappa_{s}\phi_{0,s}+J_{00}\phi_{0,t}+u_{0}s_{t+1}+\xi_{0,t+1}.$
The physical interpretation of this equation is as follows. The first term, $\eta_{t}$, is the feedforward input from the reservoir. The second term, $\kappa_{s}\phi_{0,s}$, arises from the perturbation induced in the reservoir by the auxiliary neuron 00, which is read out by neuron 00. The third term originates from self-connection of the auxiliary neuron 00.
@@
For evaluation of the second term in Eq.(A.1), we assess the order of $\chi_{ij,ts}$. We begin with $s=t-1$. Since $\chi_{ij,t\hskip 1.0ptt-1}=J_{ij}\phi^{\prime}_{j,t-1}$, we find $[\chi_{ij,t\hskip 1.0ptt-1}^{2}]\leq[J_{ij}^{2}]\sim O(1/N)$. Subsequently, for $s=t-2$, $\chi_{ij,t\hskip 1.0ptt-2}=\sum_{k}^{N}J_{ik}\phi^{\prime}_{k,t-1}\chi_{kj,t-1\hskip 1.0ptt-2}$, leading to $[\chi_{ij,t\hskip 1.0ptt-2}^{2}]\leq\sum_{k}^{N}[J_{ik}^{2}][(\chi_{kj,t-1\hskip 1.0ptt-2})^{2}]\sim O(1/N)$, where we used the result for $s=t-1$, $[(\chi_{kj,t-1\hskip 1.0ptt-2})^{2}]\sim O(1/N)$. By inductively repeating this analysis, we conclude that $[\chi_{ij,ts}^{2}]\sim O(1/N)$ for any $s<t$. Therefore, the order of the second term is evaluated as
$[\kappa_{s}^{2}]$ $=\sum_{i,j,k,l}^{N}[J_{0i}J_{j0}J_{0k}J_{l0}][\phi^{\prime}_{i,t}\phi^{\prime}_{k,t}\chi_{ij,ts}\chi_{kl,ts}]\leq\frac{g^{4}}{N^{2}}\sum_{i,j}^{N}[\chi_{ij,ts}^{2}]\sim O(1/N).$ (19)
Eventually, we obtain $[\kappa_{s}^{2}]\sim O(1/N)$ for $s<t$.
Employing Cauchy–Schwarz inequality, it is shown that all cross terms between $\eta_{t}$, $\kappa_{s}$, and $J_{00}\phi_{0,t}$ are $O(1/\sqrt{N})$.
@@
$[x_{0,t+1}^{2}]$ $=\left[\big{(}\eta_{t}+\sum_{s<t}\kappa_{s}\phi_{0,s}+J_{00}\phi_{0,t}+u_{0}s_{t+1}+\xi_{0,t+1}\big{)}^{2}\right]$ (20)
@@
Thus, to leading order, we can ignore the non-trivial terms arising from the effects of the perturbations induced by the auxiliary neuron. The statistical behavior of the auxiliary neuron is equivalent to that of the reservoir neurons, which enables the replacement of $[x_{0,t}^{2}]$ with $[x_{i,t}^{2}]$, resulting in
@@
For the scailing assumption given by Eq.(3), we can assume the inputs to be subleading, obtaining
@@
For later use, we evaluate the order of the quenched variance of $\langle x_{i}^{2}\rangle$, denoted by ${\rm{Var}}[\langle x_{i}^{2}\rangle]$, which is equivalent to the population variance because of the self-averaging property. Since the cavity method allows us to ignore the effects of perturbation induced by the auxiliary neuron, we easily calculate $[\langle x_{0}^{2}\rangle^{2}]$ as
@@
As shown in Appendix.A.2, we see $[\langle\phi_{i}\phi_{j}\rangle^{2}]_{i\neq j}\sim O(1/N)$. The self-averaging property allows us to evaluate the first term as
@@
provided that $[\langle x_{i}^{2}\rangle]\sim O(1)$.
@@ A.2 Statistics for neuronal correlations
We aim to derive the quenched average of the squared neuronal correlation, $[\langle x_{i}x_{j}\rangle^{2}]$, which is equivalent to the population average to leading order, $[\langle x_{i}x_{j}\rangle^{2}]=\frac{1}{N^{2}-N}\sum_{i,j(i\neq j)}^{N}\langle x_{i}x_{j}\rangle^{2}$. Similarly to the single neuronal statistics, the model’s symmetry ensures that $[\langle x_{i}x_{j}\rangle]=0$.
As in Appendix.A.1, we introduce two auxiliary neurons indexed by 00 and $0^{\prime}$ to the original reservoir RNN. Following the notation in Clark’s work [20], Latin indices are used to label reservoir neurons, while Greek indices denote auxiliary neurons. The dynamics of the auxiliary neurons $\mu\in\{0,0^{\prime}\}$ is described by
@@
$x_{\mu,t+1}$ $=\underbrace{\sum_{i}^{N}J_{\mu i}\phi_{i,t}}_{\equiv\eta_{\mu,t}}+\sum_{s<t}\sum_{\nu}\underbrace{\sum_{i,j}^{N}J_{\mu i}J_{j\nu}\phi^{\prime}_{i,t}\chi_{ij,ts}}_{\equiv\kappa_{\mu\nu,ts}}\phi_{\nu,s}+\sum_{\nu}J_{\mu\nu}\phi_{\nu,t}+u_{\mu}s_{t+1}+\xi_{\mu,t+1}$
$\equiv\eta_{\mu,t}+\sum_{\nu}\sum_{s<t}\kappa_{\mu\nu,ts}\phi_{\nu,s}+\sum_{\nu}J_{\mu\nu}\phi_{\nu,t}+u_{\mu}s_{t+1}+\xi_{\mu,t+1}.$
The physical interpretation of this equation is analogous to the case of single neuronal statistics (Eq.A.1): $\eta_{\mu,t}$ denotes the direct feedforward input from the reservoir to the auxiliary neuron $\mu$, and $\kappa_{\mu\nu,ts}\phi_{\nu,s}$ arises from the perturbation induced in the reservoir by the auxiliary neuron $\nu$ at time $s$, which is readout by the auxiliary neuron $\mu$ at time $t$.
Expanding $[\langle x_{0}x_{0^{\prime}}\rangle^{2}]$ using Eq.(A.2), we need to evaluate all cross terms such as $[\langle\eta_{0,t}\eta_{0^{\prime},t}\rangle^{2}]$ and $[\langle\kappa_{0\mu,ts}\kappa_{0^{\prime}\nu,ts}\rangle^{2}]$. We can accomplish this task through direct calculation, analogous to the approach used in Appendix.A.1. First,
@@
where we used in the second equation the fact that the weights connecting to auxiliary neurons are independent from the reservoir neurons’ activities. By making an ansatz, $[\langle\phi_{i}\phi_{j}\rangle^{2}]_{(i\neq j)}\sim O(1/N)$, we obtain $[\langle\eta_{0,t}\eta_{0^{\prime},t}\rangle^{2}]\sim O(1/N)$. Next, in the same manner, we obtain
$[\langle\kappa_{0\mu,ts}\kappa_{0^{\prime}\nu,ts}\rangle^{2}]$ $=\frac{g^{8}}{N^{4}}\sum_{i,j,k,l}^{N}[\langle\phi^{\prime}_{i,t}\phi^{\prime}_{k,t}\chi_{ij,ts}\chi_{kl,ts}\rangle^{2}]\leq g^{8}[\langle(\phi^{\prime}_{i,t})^{2}\rangle]^{2}[\langle\chi_{ij,ts}^{2}\rangle]^{2}\sim O(1/N^{2}).$
Repeating the same calculations, all cross terms except for the two $O(1/N)$ terms, $[\langle\eta_{0,t}\eta_{0^{\prime},t}\rangle^{2}]$ and $[\langle(u_{0}s_{t})(u_{0^{\prime}}s_{t})\rangle^{2}]=\tilde{\sigma}_{s}^{4}/N$, are shown to scale as $O(1/N^{2})$. Consequently,
@@
Replacing $[\langle x_{0}x_{0^{\prime}}\rangle^{2}]$ with $[\langle x_{i}x_{j}\rangle^{2}]_{(i\neq j)}$, we obtain

### v3

@@ VIII Discussion
The specific memory and nonlinearity requirements of a given task can be assessed using the framework proposed by Hülser et al. [38]. In this context, our results provide a practical principle for designing cost-effective RC systems: while increasing L raises computational and structural costs and yields diminishing returns for memory, it becomes essential when tasks demand higher-order nonlinear computations.
Beyond its practical usefulness, our work offers a new lens for understanding RC performance. The computational ability of RC has been predominantly discussed from the perspective of proximity to the edge of chaos [39, 40, 14, 15], typically for a fixed number of readout neurons [13, 27]. To our knowledge, this is the first study to analyze the computational ability from the novel perspective of its scaling behavior with an increasing number of readout neurons and the corresponding role of neuronal correlations. Interestingly, a recent report has suggested that the edge of chaos is not always optimal [41], but the reason for this remains unclear. The new perspective we propose may hold a crucial key to unraveling this puzzle.
It is also instructive to contrast our findings with other models of memory, such as Hopfield networks [42, 43]. In the field of Hopfield networks, it is well-established that making network activity sparse dramatically improves the capacity of associative memory [44, 45]. Since sparse activity tends to weaken correlations between neurons [46], this implies that the strength of neuronal correlations is strongly associated with the memory ability of Hopfield networks. At first glance, our work might seem to merely replicate these prior findings. However, there is a fundamental difference between them. The mechanism by which Hopfield networks store information utilizes the fixed points of attractor dynamics, which can be described as static and long-term memory. RC, in contrast, takes the approach of temporarily retaining time-series information through reverberating neural activity within a reservoir network, which is a form of dynamic and short-term memory (or working memory). Given that their memory mechanisms are entirely different, there is no guarantee that a principle governing static memory would apply equally to dynamic memory.
@@
Our findings have implications that extend beyond the framework of RC. First, in neuroscience, the mammalian cortex in the awake state typically operates in an asynchronous regime characterized by near-zero neuronal correlations [47, 48, 49], while recent studies have shown the level of correlation is structured [50, 51] and influenced by factors such as stimulus drive [52, 53] and attention [54, 55]. The functional impacts of the neuronal correlations have been extensively investigated through the lens of neural manifolds [56, 57, 58]. Within this framework, a neuronal network with a high level of correlated neuronal activity occupies a restricted region of $N$-dimensional phase space, the linear dimension of which is quantified by the effective dimension [59]. In the context of population coding, a well-established trade-off exists between representational richness and robustness: high-dimensional (i.e., low-correlation) activity enables complex and diverse information encoding, while low-dimensional (i.e., high-correlation) activity supports redundant encoding and noise robustness [57, 58]. Existing theoretical frameworks, however, have primarily focused on the implications of neuronal correlations for static information processing, i.e., mapping static inputs to outputs. Our findings suggest a novel role for neuronal correlations in dynamic information processing, offering a new hypothesis regarding their functional significance in the brain.
@@
Data availability: The source code used to generate all figures, except for Fig.8, is available at [65].
@@ Appendix A THEORETICAL ANALYSIS ON STATISTICS OF NEURONAL ACTIVITY
In this section, we derive statistics of neuronal activity based on dynamical cavity approach [26, 66]. Several results and methodologies introduced here are frequently utilized in deriving memory capacity in Appendix B.
In deriving single neuronal statistics such as the variance of $x_{i}(t)$ (see Appendix A.1) in the limit of $N\to\infty$, the same result can be obtained via dynamical mean-field theory [18, 25] . However, since the dynamical mean-field theory transforms the $N$-body system into the effective single-body system, it is inherently unable to evaluate many-body neuronal statistics, such as neuronal correlations (see Appendix A.2) and memory capacity(see Appendix B.1). In contrast, the dynamical cavity method is, in principle, applicable to $n$-body neuronal statistics for any $n$. Furthermore, while the aforementioned limitation of dynamical mean-field theory has recently been addressed through subleading corrections to the saddle-point solution [67, 68], the dynamical cavity method remains advantageous for its conceptual clarity.
@@
Before discussing the details, we first outline the necessity of employing the dynamical cavity method. Consider calculating the quenched variance of neuronal activity, $x_{i,t}^{2}]$. For simplicity, we ignore input and noise signals. According to time evolution equation of neurons (Eq.(1)), $x_{i,t}$ is the sum of a large number ($N\gg 1$) of mutually independent quantities, $J_{ij}\phi(x_{j,t-1})$. This implies that $x_{i,t}$ follows a Gaussian distribution due to the central limit theorem, and thus, it is sufficient to determine the complete statistics of $x_{i,t}$ by calculating its mean and variance. Given the symmetry of the model, we have $x_{i,t}^{2}]=0$. On the other hand, we can intuitively calculate $x_{i,t}^{2}]$ as:
$[x_{i,t}^{2}]=\sum_{j,k}^{N}\left[J_{ij}J_{ik}\phi_{j,t-1}\phi_{k,t-1}\right]\overset{?}{=}\sum_{j,k}^{N}\left[J_{ij}J_{ik}\right]\left[\phi_{j,t-1}\phi_{k,t-1}\right]=g^{2}\left[\phi_{j,t-1}^{2}\right]=g^{2}\langle\phi(x)^{2}\rangle_{x\sim\mathcal{N}(0,[x_{i,t-1}^{2}])},$ (16)
yielding time series of $x_{i,t}^{2}]$. Here we assumed in the second equation that the recurrent weights, $J_{ij}$ and $J_{ik}$, are independent of the neuronal activities, $\phi_{j}$ and $\phi_{k}$. Fortunately, this assumption is validated in the limit of $N\to\infty$, as shown with the dynamical mean-field theory. However, this assumption is likely to break down when finite size effects are taken into account. To circumvent this challenge, we need to employ the dynamical cavity method.
@@ A.1 Single neuronal statistics
We aim to derive the quenched average of the time-variance of neuronal activity, $\langle x_{i}^{2}\rangle]$. Note that due to the self-averaging, it is equivalent to the population average of time-variance to leading order, i.e. $\langle x_{i}^{2}\rangle]=\frac{1}{N}\sum_{i}^{N}\langle x_{i}^{2}\rangle$. The symmetry of the model setting ensures that the quenched average of the time-mean, $\langle x_{i}\rangle]$, vanishes.
The overview of dynamical cavity method is described in Fig.8. One auxiliary neuron, indexed by 0, is added to the original reservoir RNN whose neurons are indexed by $i=1,\cdots,N$. The original neuron is perturbated by the addition of the auxiliary neuron. By denoting the perturbation applied to the $i$th neuron at time $t$ by $\delta_{i,t}$, we describe the dynamics of the auxiliary neuron as
@@
We define $\chi_{ij,t\ t-\tau}\equiv\delta x_{i,t}/\delta I_{j,t-\tau}$ as the linear susceptibility of the $i$th neuron at time t, induced by an infinitesimal external inputs applied to the $j$th neuron at time $t-\tau\ (\tau\geq 1)$, $I_{j,t-\tau}$. Then, as the perturbation to $i$th neuron at time $t$, $\delta_{i,t}$, is the sum of all direct perturbations to $j\in\{1,\cdots,N\}$ at time $t-\tau\ (\tau\geq 1)$ weighted by the corresponding susceptibility $\chi_{ij,t\ t-\tau}$, we obtain
$\delta_{i,t}=\sum_{\tau\geq 1}\sum_{j=1}^{N}\chi_{ij,t\ t-\tau}J_{j0}\phi_{0,t-\tau-1}.$ (18)
@@
$x_{0,t+1}$ $=\underbrace{\sum_{i=1}^{N}J_{0i}\phi_{i,t}}_{\equiv\eta_{t}}+\sum_{\tau\geq 1}\underbrace{\sum_{i,j}^{N}\chi_{ij,t\ t-\tau}J_{0i}J_{j0}\phi^{\prime}_{i,t}}_{\equiv\kappa_{\tau}}\phi_{0,t-\tau-1}+J_{00}\phi_{0,t}+u_{0}s_{t+1}+\xi_{0,t+1}$
$\equiv\eta_{t}+\sum_{\tau\geq 1}\kappa_{\tau}\phi_{0,t-\tau-1}+J_{00}\phi_{0,t}+u_{0}s_{t+1}+\xi_{0,t+1}.$
The physical interpretation of this equation is as follows. The first term, $\eta_{t}$, is the feedforward input from the reservoir. The second term, $\kappa_{\tau}\phi_{0,t-\tau-1}$, arises from the perturbation induced in the reservoir by the auxiliary neuron 0, which is read out by neuron 0. The third term originates from self-connection of the auxiliary neuron 0.
@@
For evaluation of the second term in Eq.(A.1), we assess the order of $\chi_{ij,t\ t-\tau}$. We begin with $\tau=1$. Since $\chi_{ij,t\hskip 1.0ptt-1}=J_{ij}\phi^{\prime}_{j,t-1}$, we find $\chi_{ij,t\hskip 1.0ptt-1}^{2}]\leq[J_{ij}^{2}]\sim O(1/N)$. Subsequently, for $\tau=2$, $\chi_{ij,t\hskip 1.0ptt-2}=\sum_{k}^{N}J_{ik}\phi^{\prime}_{k,t-1}\chi_{kj,t-1\hskip 1.0ptt-2}$, leading to $\chi_{ij,t\hskip 1.0ptt-2}^{2}]\leq\sum_{k}^{N}[J_{ik}^{2}][(\chi_{kj,t-1\hskip 1.0ptt-2})^{2}]\sim O(1/N)$, where we used the result for $\tau=1$, $(\chi_{kj,t-1\hskip 1.0ptt-2})^{2}]\sim O(1/N)$. By inductively repeating this analysis, we conclude that $\chi_{ij,t\ t-\tau}^{2}]\sim O(1/N)$ for any $\tau\geq 1$. Therefore, the order of the second term is evaluated as
$[\kappa_{\tau}^{2}]$ $=\sum_{i,j,k,l}^{N}[J_{0i}J_{j0}J_{0k}J_{l0}][\phi^{\prime}_{i,t}\phi^{\prime}_{k,t}\chi_{ij,t\ t-\tau}\chi_{kl,t\ t-\tau}]\leq\frac{g^{4}}{N^{2}}\sum_{i,j}^{N}[\chi_{ij,t\ t-\tau}^{2}]\sim O(1/N).$ (22)
Eventually, we obtain $\kappa_{\tau}^{2}]\sim O(1/N)$ for $\tau\geq 1$.
Employing Cauchy–Schwarz inequality, it is shown that all cross terms between $\eta_{t}$, $\kappa_{\tau}$, and $J_{00}\phi_{0,t}$ are $O(1/\sqrt{N})$.
@@
$[x_{0,t+1}^{2}]$ $=\left[\big{(}\eta_{t}+\sum_{\tau\geq 1}\kappa_{\tau}\phi_{0,t-\tau-1}+J_{00}\phi_{0,t}+u_{0}s_{t+1}+\xi_{0,t+1}\big{)}^{2}\right]$ (23)
@@
Thus, to leading order, we can ignore the non-trivial terms arising from the effects of the perturbations induced by the auxiliary neuron. The statistical behavior of the auxiliary neuron is equivalent to that of the reservoir neurons, which enables the replacement of $x_{0,t}^{2}]$ with $x_{i,t}^{2}]$, resulting in
@@
For the scaling assumption given by Eq.(4), we can assume the inputs to be subleading, obtaining
@@
For later use, we evaluate the order of the quenched variance of $\langle x_{i}^{2}\rangle$, denoted by ${\rm{Var}}[\langle x_{i}^{2}\rangle]$, which is equivalent to the population variance because of the self-averaging property. Since the cavity method allows us to ignore the effects of perturbation induced by the auxiliary neuron, we easily calculate $\langle x_{0}^{2}\rangle^{2}]$ as
@@
As shown in Appendix A.2, we see $\langle\phi_{i}\phi_{j}\rangle^{2}]_{i\neq j}\sim O(1/N)$. The self-averaging property allows us to evaluate the first term as
@@
provided that $\langle x_{i}^{2}\rangle]\sim O(1)$.
@@ A.2 Statistics for neuronal correlations
We aim to derive the quenched average of the squared neuronal correlation, $\langle x_{i}x_{j}\rangle^{2}]$, which is equivalent to the population average to leading order, $\langle x_{i}x_{j}\rangle^{2}]=\frac{1}{N^{2}-N}\sum_{i,j(i\neq j)}^{N}\langle x_{i}x_{j}\rangle^{2}$. Similarly to the single neuronal statistics, the model’s symmetry ensures that $\langle x_{i}x_{j}\rangle]=0$.
As in Appendix A.1, we introduce two auxiliary neurons indexed by 0 and $0^{\prime}$ to the original reservoir RNN. Following the notation in Clark’s work [26], Latin indices are used to label reservoir neurons, while Greek indices denote auxiliary neurons. The dynamics of the auxiliary neurons $\mu\in\{0,0^{\prime}\}$ is described by
@@
$x_{\mu,t+1}$ $=\underbrace{\sum_{i}^{N}J_{\mu i}\phi_{i,t}}_{\equiv\eta_{\mu,t}}+\sum_{\tau\geq 1}\sum_{\nu}\underbrace{\sum_{i,j}^{N}J_{\mu i}J_{j\nu}\phi^{\prime}_{i,t}\chi_{ij,t\ t-\tau}}_{\equiv\kappa_{\mu\nu,t\ t-\tau}}\phi_{\nu,t-\tau-1}+\sum_{\nu}J_{\mu\nu}\phi_{\nu,t}+u_{\mu}s_{t+1}+\xi_{\mu,t+1}$
$\equiv\eta_{\mu,t}+\sum_{\nu}\sum_{\tau\geq 1}\kappa_{\mu\nu,t\ t-\tau}\phi_{\nu,t-\tau-1}+\sum_{\nu}J_{\mu\nu}\phi_{\nu,t}+u_{\mu}s_{t+1}+\xi_{\mu,t+1}.$
The physical interpretation of this equation is analogous to the case of single neuronal statistics (Eq.A.1): $\eta_{\mu,t}$ denotes the direct feedforward input from the reservoir to the auxiliary neuron $\mu$, and $\kappa_{\mu\nu,t\ t-\tau}\phi_{\nu,t-\tau-1}$ arises from the perturbation induced in the reservoir by the auxiliary neuron $\nu$ at time $t-\tau-1$, which is read out by the auxiliary neuron $\mu$ at time $t$.
Expanding $\langle x_{0}x_{0^{\prime}}\rangle^{2}]$ using Eq.(A.2), we need to evaluate all cross terms. We can accomplish this task through direct calculation, analogous to the approach used in Appendix A.1. First,
@@
where we used in the second equation the fact that the weights connecting to auxiliary neurons are independent from the reservoir neurons’ activities. By making an ansatz, $\langle\phi_{i}\phi_{j}\rangle^{2}]_{(i\neq j)}\sim O(1/N)$, we obtain $\langle\eta_{0,t}\eta_{0^{\prime},t}\rangle^{2}]\sim O(1/N)$.
Next, calculating the cross term between the second and third terms yields
$\left[\left\langle\sum_{\mu}\sum_{\tau\geq 1}\kappa_{0\mu,t\ t-\tau}\phi_{\mu,t-\tau-1}\sum_{\nu}J_{0^{\prime}\nu}\phi_{\nu,t}\right\rangle^{2}\right]$ $=$ $\frac{g^{2}}{N}\sum_{\mu,\nu,\rho}\sum_{\tau,\tau^{\prime}}\left[\langle\kappa_{0\mu,t\ t-\tau}\phi_{\mu,t-\tau-1}\phi_{\nu,t}\rangle\langle\kappa_{0\rho,t\ t-\tau}\phi_{\rho,t-\tau^{\prime}-1}\phi_{\nu,t}\rangle\right]$
$=$ $\frac{g^{2}}{N}\sum_{\mu,\nu,\rho}\sum_{\tau,\tau^{\prime}}\sum_{i,j,k,l}^{N}[J_{0i}J_{j\mu}J_{0k}J_{l\nu}][\langle\phi^{\prime}_{i,t}\chi_{ij,t\ t-\tau}\phi_{\mu,t-\tau-1}\phi_{\nu,t}\rangle$
$\hskip 120.0pt\times\langle\phi^{\prime}_{k,t}\chi_{kl,t\ t-\tau^{\prime}}\phi_{\rho,t-\tau^{\prime}-1}\phi_{\nu,t}\rangle]$
$=$ $\frac{g^{6}}{N^{3}}\sum_{\mu}\sum_{\tau,\tau^{\prime}}\sum_{i,j}^{N}[\langle\phi^{\prime}_{i,t}\rangle^{2}\langle\chi_{ij,t\ t-\tau}\chi_{ij,t\ t-\tau^{\prime}}\rangle\langle\phi_{\mu,t-\tau-1}\phi_{\mu,t}\rangle\langle\phi_{\mu,t-\tau^{\prime}-1}\phi_{\mu,t}\rangle]$
$\sim$ $O(1/N^{2}),$ (37)
where we used $\chi_{ij,t\ t-\tau}^{2}\sim O(1/\sqrt{N})$ for any $\tau\geq 1$, as shown before. Therefore, this cross term becomes sub-leading in $\langle x_{0}x_{0^{\prime}}\rangle^{2}]$.
The cross term between the first and second terms vanishes as
$\left[\left\langle\eta_{0,t}\sum_{\mu}\sum_{\tau\geq 1}\kappa_{0^{\prime}\mu,t\ t-\tau}\phi_{\mu,t-\tau-1}\right\rangle^{2}\right]$ $=$ $\left[\left(\sum_{i,j,k}^{N}\sum_{\mu}\sum_{\tau\geq 1}J_{0i}J_{0^{\prime}j}J_{k\mu}\left\langle\phi_{i,t}\phi^{\prime}_{j,t}\phi_{\mu,t-\tau-1}\right\rangle\right)^{2}\right]$
$=$ $\left[\left(\sum_{i,j,k}^{N}\sum_{\mu}\sum_{\tau\geq 1}J_{0i}J_{0^{\prime}j}J_{k\mu}\left\langle\phi_{i,t}\phi^{\prime}_{j,t}\right\rangle\left\langle\phi_{\mu,t}\right\rangle\right)^{2}\right]$
$=$ $0,$
where the second equality holds because the activity of the reservoir neuron is independent of that of the auxiliary neurons.
Repeating these calculations, all cross terms are shown to scale as $O(1/N^{2})$, except for the two $O(1/N)$ terms, $\langle\eta_{0,t}\eta_{0^{\prime},t}\rangle^{2}]$ and $\langle(u_{0}s_{t})(u_{0^{\prime}}s_{t})\rangle^{2}]=\tilde{\sigma}_{s}^{4}/N$. Consequently,
@@
Replacing $\langle x_{0}x_{0^{\prime}}\rangle^{2}]$ with $\langle x_{i}x_{j}\rangle^{2}]_{(i\neq j)}$, we obtain